"One number is four times another number. The greater of the two numbers is 20 less than three times the difference between the two numbers. Find the numbers." There are two possible answers. Do these equations make sense: 4a=b and b-20=3(b-a)?
Answer by mananth(16946)

4a=b and
b=3(b-a)-20
4a=3(4a-a)-20
4a=3*3a-20
4a-9a=-20
-5a=-20
/-5
a=4
b=16
..
m.ananth@hotmail.ca